OpenStudy (anonymous):

what is the simplified form of 2/x^2-x-1/x

OpenStudy (anonymous):

options are:

OpenStudy (anonymous):

I think it's C but not sure. Anyone out there that could help?

OpenStudy (anonymous):

@zepdrix hiii you busy rn hun?

zepdrix (zepdrix):

\[\large\rm \frac{2}{x^2}-x-\frac{1}{x}\]This?

OpenStudy (anonymous):

no, lemme rephrase it: \[\frac{ 2 }{ x^2-x } - \frac{ 1 }{ x }\]

OpenStudy (anonymous):

@zepdrix like that ^^

zepdrix (zepdrix):

Oh, learn to format it correctly you silly billy :U 2/(x^2-x)-1/x

OpenStudy (anonymous):

sorryy haha. and yea thats it

zepdrix (zepdrix):

3-x in top? Mmmm ya I think that's right! :) You factor the first fraction into x(x-1) revealing that the second fraction needs an (x-1) top and bottom. and then boom bam, you end up with 3-x after combining some stuff right?

OpenStudy (anonymous):

yes. so it's really C? awesome! could i ask one more question? I'll tag u in it